OpenStudy (anonymous):

WILL GIVE YOU FAN AND MEDAL!!!!!! Part A: The area of a square is (9x2 - 12x + 4) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. (3 points) Part B: The area of a rectangle is (25x2 - 16y2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work. (2 points) Part C: The volume of a rectangular box is (x3 - 7x2 - 9x + 63) cubic units. Determine the dimensions of the rectangular box by factoring the volume expression completely. Show your work. (5 points)

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Do you know what the formula for the area of a square is?

OpenStudy (igreen):

The formula is: \(A = s^2\) Plug in the area: \((9x^2 - 12x + 4) = s^2\) To solve you have to square both sides: \(\sqrt{(9x^2 - 12x + 4)} = s\) Which simplifies to: \(\sqrt{(2-3x)^2}\)

OpenStudy (anonymous):

so sorry was afk

OpenStudy (igreen):

No problem..I think that solves Part A for you..

OpenStudy (anonymous):

so all that is A right?

OpenStudy (anonymous):

ok, ty on to B i see it and understand

OpenStudy (igreen):

The formula for the area of a rectangles is \(A = lw\) Area = \((25x^2 - 16y^2)\) If we factor this into two separate expressions, they will be the length and the width. If you factor this you get \((5x - 4y)(5x+4y)\), which represents the dimensions of the rectangle.

OpenStudy (anonymous):

and this is all of B right?

OpenStudy (igreen):

Yep.

OpenStudy (anonymous):

awesome on to C :)

OpenStudy (igreen):

\( (x^3 - 7x^2 - 9x + 63)\) For part C, we have to factor this too, but into three parts instead of two, because the area of a rectangular prism is \(A = lwh\). We get \((x-7)(x-3)(x+3)\) when we factor, so those will be the dimensions of the rectangular box.

OpenStudy (igreen):

That takes care of C..